What is the strength of the electric field between two parallel conducting plates separated by 1.00 cm and having a potential difference (voltage) between them of 1.50×10⁴V?

1
Expert's answer
2022-02-15T09:57:54-0500

Given:

"d=\rm 1.00\: cm=0.010 \:m"

"V=1.50×10^4\:\rm V"


The strength of the electric field between two parallel conducting plates

"E=\frac{V}{d}=\rm \frac{1.50×10^4\:\rm V}{0.010\: m}=1.5*10^6\:\rm V/m"